The base BC of triangle ABC is trisected at X and Y .
Show that |AX|
2 + |AY|
2 + 4|XY|
2 = |AB|
2 + |AC|
2
1
Expert's answer
2019-10-03T08:44:37-0400

Let "BX=\frac{1}{3}BC" and "BY=\frac{2}{3}BC" .

Consider "\overrightarrow{AB}" and "\overrightarrow{AC}" as a basis of a vector space.

We have

1)"\overrightarrow{BC}=\overrightarrow{AC}-\overrightarrow{AB}" , so "|\overrightarrow{BC}|^2=|\overrightarrow{AC}|^2+|\overrightarrow{AB}|^2-2(\overrightarrow{AC},\overrightarrow{AB})." That is "(\overrightarrow{AC},\overrightarrow{AB})=\frac{1}{2}(AC^2+AB^2-BC^2)"

2)"\overrightarrow{AX}=\overrightarrow{AB}+\overrightarrow{BX}=\overrightarrow{AB}+\frac{1}{3}\overrightarrow{BC}=\overrightarrow{AB}+\frac{1}{3}(\overrightarrow{AC}-\overrightarrow{AB})="

"=\frac{2}{3}\overrightarrow{AB}+\frac{1}{3}\overrightarrow{AC}", so "|\overrightarrow{AX}|^2=\frac{4}{9}|\overrightarrow{AB}|^2+\frac{1}{9}|\overrightarrow{AC}|^2+\frac{4}{9}(\overrightarrow{AB},\overrightarrow{AC})="

"=\frac{4}{9}AB^2+\frac{1}{9}AC^2+\frac{2}{9}(AC^2+AB^2-BC^2)="

"=\frac{2}{3}AB^2+\frac{1}{3}AC^2-\frac{2}{9}BC^2"

3)"\overrightarrow{AY}=\overrightarrow{AB}+\overrightarrow{BY}=\overrightarrow{AB}+\frac{2}{3}\overrightarrow{BC}=\overrightarrow{AB}+\frac{2}{3}(\overrightarrow{AC}-\overrightarrow{AB})="

"=\frac{1}{3}\overrightarrow{AB}+\frac{2}{3}\overrightarrow{AC}", so "|\overrightarrow{AY}|^2=\frac{1}{9}|\overrightarrow{AB}|^2+\frac{4}{9}|\overrightarrow{AC}|^2+\frac{4}{9}(\overrightarrow{AB},\overrightarrow{AC})="

"\frac{1}{9}AB^2+\frac{4}{9}AC^2+\frac{2}{9}(AC^2+AB^2-BC^2)="

"=\frac{1}{3}AB^2+\frac{2}{3}AC^2-\frac{2}{9}BC^2"

4)From 2 and 3 we obtain "AX^2+AY^2=\bigl(\frac{2}{3}AB^2+\frac{1}{3}AC^2-\frac{2}{9}BC^2\bigr)+"

"+\bigl(\frac{1}{3}AB^2+\frac{2}{3}AC^2-\frac{2}{9}BC^2\bigr)=AB^2+AC^2-\frac{4}{9}BC^2"

Since "XY=\frac{1}{3}BC", we have "XY^2=\frac{1}{9}BC^2" , so "AX^2+AY^2=AB^2+AC^2-\frac{4}{9}BC^2=AB^2+AC^2-4XY^2" . Moving "4XY^2" to the left side, we obtain "AX^2+AY^2+4XY^2=AB^2+AC^2" .